LIE SUPERALGEBRAS GRADED BY THE ROOT SYSTEM A(m, n)

GEORGIA BENKART ALBERTO ELDUQUE

Abstract. We determine the Lie superalgebras that are graded by the root systems of the basic classical simple Lie superalgebras of type A(m, n).